OpenStudy (anonymous):

On the way to work Chad averaged 70 mph and on the return trip home he averaged 60 mph. It took Chad (1/4) of an hour longer to get home because of the heavier traffic. How far does Chad travel to work?*

OpenStudy (anonymous):

The distance to work is R*T or 70t, where t is the time to get to work The distance from work is R*T or 60(t+.25) Since these distances are the same we may equate them: 70t=60(t+.25) Now solve this for t. (cont)

OpenStudy (anonymous):

70t=60t+15 10t=15 t=3/2=1.5 Now you have to sub this into one of the equations to get the distance: distance = 70(1.5) = 105 or distance = 60(1.5+.25) = 105

OpenStudy (anonymous):

Holy Giabroni, 105 miles to work!